A certain number of the Durban car commuters has a sample mean time of 35.8 to commute to work daily. Assuming that the population standard deviation is 11 minutes, standard error is 1.1 minutes and that commuting times are normally distributed. Identify the size of the sample.

Select one:

a. 70

b. 20

c. 100

d. 60


1
Expert's answer
2021-08-29T17:41:00-0400
"SE=\dfrac{\sigma}{\sqrt{n}}"

"n=(\dfrac{\sigma}{SE})^2"

"n=(\dfrac{11}{1.1})^2=100"

c. 100